OpenStudy (anonymous):

Using the substitution method, solve the equations for x and y. -x-5y=-2 2x-7y=25

OpenStudy (ash2326):

1st......-x-5y=-2 2nd......2x-7y=25 find x from 1st -x-5y=-2 -5y+2=x x=2-5y substitute thin in second 2(2-5y)-7y=-2 4-10y-7y=-2 -17y=-6 y=6/17 x=2-5y x=2-5*6/17 x=2-30/17 x=(34-30)/17 x=4/17 y=6/17